# Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Finetuning

### v2

@@ Abstract
Reinforcement finetuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely. Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear. We propose Arbitrary Entropy Policy Optimization (AEPO), which eliminates entropy collapse by replacing entropy bonuses with REINFORCE policy gradient on temperature-adjusted distributions and stabilizing entropy through temperature regulation. AEPO integrates three key designs: policy gradient as regularization, distribution as regularization, and REINFORCE as regularization, enabling precise entropy control without distorting optimization. Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers.
@@ 1 Introduction
Reinforcement finetuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024). Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025). However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (Yu et al., 2025; Li et al., 2025a; Zhang et al., 2025). This severely restricts the ability of RFT to discover diverse reasoning strategies. Existing remedies typically rely on entropy regularization and clip refine, which inevitably introduce bias and instability into the optimization objective, creating a trade-off between exploration and stability (Hou et al., 2025; Cui et al., 2025; Cheng et al., 2025; Shen, 2025). EFRame is the first to achieve large-scale entropy control, but it remains a qualitative analysis without conducting a quantitative study of entropy (Wang et al., 2025a).
@@ 4.2 AEPO
Benchmarks AIME24 AMC Challenge math GSM8K MATH Minerva math Olympiad Average
@@ 5.1 Experimental setup
Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Zhong et al., 2023), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023). These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
@@ 5.3 Entropy and exploration
AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg
@@
AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg

### v4

# Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Fine-tuning

@@ Abstract
Reinforcement fine-tuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely. Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear. We propose Arbitrary Entropy Policy Optimization (AEPO), which eliminates entropy collapse by replacing entropy bonuses with REINFORCE policy gradient on temperature-adjusted distributions and stabilizing entropy through temperature regulation. AEPO integrates three key designs: policy gradient as regularization, distribution as regularization, and REINFORCE as regularization, enabling precise entropy control without distorting optimization. Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers. Our code is available at https://github.com/597358816/AEPO.
@@ 1 Introduction
Reinforcement fine-tuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024). Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025). However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (Yu et al., 2025; Li et al., 2025a; Zhang et al., 2025). This severely restricts the ability of RFT to discover diverse reasoning strategies. Existing remedies typically rely on entropy regularization and clip refine, which inevitably introduce bias and instability into the optimization objective, creating a trade-off between exploration and stability (Hou et al., 2025; Cui et al., 2025; Cheng et al., 2025; Shen, 2025). EFRame is the first to achieve large-scale entropy control, but it remains a qualitative analysis without conducting a quantitative study of entropy (Wang et al., 2025a).
@@ 4.2 AEPO
Benchmarks AIME24 AMC College Math GSM8K MATH Minerva Math Olympiad Average
@@ 5.1 Experimental setup
Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Tang et al., 2024), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023). These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
@@ 5.3 Entropy and exploration
AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
@@
AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg